OpenStudy (anonymous):

Choose the equation of the line passing through the point (6, -2) and parallel to y = x + 7. y = x + 8 y = x - 4 y = x + 4 y = x - 8

jimthompson5910 (jim_thompson5910):

each answer choice has the same slope (which is 1), so we don't have to worry about the slope

jimthompson5910 (jim_thompson5910):

all we need to do is plug in the point (6,-2) into each answer choice and see if we get a true equation or not

jimthompson5910 (jim_thompson5910):

for instance, in choice A y = x + 8 -2 = 6 + 8 ... plug in (6,-2) -2 = 14 that's false, so (6,-2) is NOT on the line y = x + 8 so choice A is out

jimthompson5910 (jim_thompson5910):

see how to do this?

OpenStudy (anonymous):

Yes thank you I think I can do it now.

jimthompson5910 (jim_thompson5910):

alright great, yw

OpenStudy (whpalmer4):

You could also construct the equation: we know the slope is 1, and a line with slope 1 going through that point would be\[y- (-2) = (1)(x-6)\]\[y+2=x-6\]\[y+2-2=x-6+2\]\(you can finish the simplification)

OpenStudy (whpalmer4):

(I used the point-slope formula for a line of slope \(m\) going through point \((x_0,y_0)\):\[y-y_0 = m(x-x_0)\]

OpenStudy (whpalmer4):

when I was learning this stuff, we never had multiple-choice questions, we always had to figure the darn thing out ourselves :-)

OpenStudy (anonymous):

Thanks whpalmer4 I think I understood it more now. Thank You!!

jimthompson5910 (jim_thompson5910):

whpalmer4's way is probably better since you won't have multiple choice to rely on all the time but if you do have multiple choice answers, you can use this quick method to check

OpenStudy (anonymous):

Okay jim_thompson5910 thanks!

OpenStudy (whpalmer4):

Yep. I'm not generally in favor of multiple-choice exams, but if they're going to give them to you, you might as well take advantage of the opportunities to solve or check your work in any way available!

OpenStudy (anonymous):

Ya so I think the answer is the last one right?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (whpalmer4):

uh, let's see: point is (6,-2) right? y = x-8 -2 = 6-8 -2 = -2 looks good.

OpenStudy (whpalmer4):

your medal is in the mail :-)

OpenStudy (anonymous):

Thank you guys! You really helped me a lot since I had to do this for an exam. And thank you whpalmer4 for the medal!!

OpenStudy (anonymous):

y = x - 4

OpenStudy (whpalmer4):

@dpasingh you think y=x-4 goes through the point (6,-2)? Let's check! y = x-4 (-2) = (6)-4 -2 = 2 Are you sure about that?